if amrik is three times chantals age, and in 4 years will be twice her age, how old are they now
Answer by mananth(16946)

chantal's age =x
Amrik's age = 3x
ages after 4 years the ages will be
Chantal= x+4
Amrik = 3x+4
(3x+4) = 2(x+4)
3x+4 = 2x+8
3x-2x=8-4
x= 4 -------- chantal's age
Amrik's age = 3x= 3*4 =12 years